HUBBLE SPACE TELESCOPE DAILY REPORT #842 PERIOD COVERED: 0800 EST Daily Status Report as of 071/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 3591 (Massive Stars in Starburst Galaxies - FOC Images) The Faint Object Camera (f/96) was used to make observations of the star bursting galaxy IZW18 to investigate the physics of its massive stars, and the processes which drive the galaxy's formation and evolution. The observations were executed and scheduled, and no problems were reported. 1.2 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.3 Completed FOC/96 4800 (Stellar Content of Wolf-Rayet Galaxies) The Faint Object Camera (f/96) was used to make one image of the Wolf-Rayet Galaxy MARK1236. The observation was reported to have produced an excellent image of the galaxy. 1.4 Completed HSP/POL 3557 (Polarimetric Observations of the Crab Nebula Pulsar in the UV) The High Speed Photometer was used to make ultraviolet polarimetric observations of the Crab Nebula Pulsar. Simultaneous radio telescope observations will be made, and a comparison of the pulse profiles and polarizations will be made. The onboard target acquisitions were successful. Four polarimetric observations were made. One of the observations was not received, because of a rejected downlink. The other three observations were reported to be okay. 1.5 Completed WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to make four low latitude observations which were taken in parallel with the prime science observations. All four observations were reported to be satisfactory. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 9 Successful : 9 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: HSTAR 3919 - The EPS/DMS group is investigating a telemetry channel failure on DIU2. Seven monitors were reported to have gone to their zero state (i.e., they showed miscompares) at Friday (071/0002Z). Preliminary data indicates possible failure of -AA Trim Relay at the same time. 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of FOS/RD 4698 (FOS Dark Monitoring) 3.2 Complete Two Sets of HSP/POL 3557 (Polarimetric Observations of the Crab Pulsar in the UV) 3.3 Complete Two Sets of WFC 1306 (GTO Low Latitude Parallel Program) 3.4 Complete PC 4347 (WF/PC Calibration: Flood Retention Test) 3.5 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete FOC 4800 (Stellar Content of Wolf-Rayet Galaxies) 3.8 Complete GHRS/2 2536 (Deuterium in the Local Interstellar Gas) 3.9 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.10 Complete FOS/RD 2563 (Supernova Intensive Study) 3.12 Complete WFC 3589 (Parallel High Resolution Imaging of Defuse Objects) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-